OpenStudy (anonymous):

The variables x and y are related by the equation y = (1- 2x)^3 . Find the rate of change of x at the instant when x = 0.3, given that y is changing at a rate of 2 units per second at this instant.

OpenStudy (amistre64):

take the derivative ....

OpenStudy (amistre64):

y = (1- 2x)^3 y' = 3(1- 2x)^2 (-2x')

OpenStudy (amistre64):

they tell you y' = 2 and x = 0.3 ... find x'

OpenStudy (anonymous):

ok